OpenStudy (anonymous):

In the equation x^2+y^2-9x+8y=5, at what points is the the tangent line horizontal? vertical? The curve has a ___horizontal or vertical____ tangent line when x=_____. The curve has a __horizontal or vertical____ tangent line when y=______. Please explain:) thank you!! and also, dy/dx of the equation is (9-2x)/(2y+8) :)

terenzreignz (terenzreignz):

It all boils down to what you know about the derivative dy/dx and how it relates to the tangent line. dy/dx is the *slope* of the tangent line, yes?

OpenStudy (anonymous):

yes:) 

terenzreignz (terenzreignz):

Simple. A horizontal line has a zero-slope. A vertical line has an undefined slope. So... for what values of x and y would the slope dy/dx be zero?

OpenStudy (anonymous):

how would i find that? 9-2x=0 solve for x? and 2y+8=0 solve for y?

terenzreignz (terenzreignz):

Don't touch the denominator yet, we're still on the horizontal tangent line. To make a fraction zero, it suffices to make the numerator equal to zero. So yes, 9 - 2x = 0 and solve for x will give you the x-values that give a horizontal tangent line. On the flipside, an undefined tangent line occurs when the denominator is zero, so to get the values that give a vertical tangent line, solve 2y + 8 = 0 and solve. ^_^

OpenStudy (anonymous):

ohh okay... so right now, just solving for x? so -2x=9 so x = -2/9?

terenzreignz (terenzreignz):

oops... try again

OpenStudy (anonymous):

oh oops haha x= -9/2 ?

terenzreignz (terenzreignz):

no. Take a deep breath and solve 9 - 2x = 0 again...

OpenStudy (anonymous):

ohh okay... umm  lol oops... my bad :P